OpenStudy (danieke):

The price of a gallon of unleaded gas was $2.82 yesterday. Today, the price rose to $2.89. Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (sasquat):

All u gotta do to this one is 2.89-2.82= 0.07 0.07 x 100 = 7%

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

and no

OpenStudy (sasquat):

whatta ya mean no

OpenStudy (misty1212):

while \[.07=7\%\] that just means 7 cents is 7 percent of a dollar

OpenStudy (misty1212):

once you know the increase is 7 cents, you need to compare to the original amount compute \[\frac{7}{282}\] and turn that number in to a percent

OpenStudy (sasquat):

whered u get 282 from

OpenStudy (rocketisawesome):

Weirdly I did this exact question about two weeks ago so I can help you.

OpenStudy (sasquat):

no thats wrong cuz 2.89 is the 100%, its what u take 2.82 out of

OpenStudy (rocketisawesome):

2.12%

OpenStudy (rocketisawesome):

thats the answer

OpenStudy (sasquat):

how did u do that tho

OpenStudy (rocketisawesome):

2.12% percent

OpenStudy (rocketisawesome):

I did it like 2 weeks ago literally I got help with a tutor

OpenStudy (rocketisawesome):

we did the question together and got 2.12 percent

OpenStudy (sasquat):

k but how????

OpenStudy (misty1212):

the answer is what you get when you divide \[7\div282\] and turn it in to a percent

OpenStudy (rocketisawesome):

hold on

OpenStudy (sasquat):

282 doesnt make sense

OpenStudy (rocketisawesome):

let me look back in my snapshots

OpenStudy (rocketisawesome):

she clearly means 2.82

OpenStudy (rocketisawesome):

@danieke the answer is 2.12 percent

OpenStudy (sasquat):

yeah but how is the answer 2.12 can u show ur work

OpenStudy (misty1212):

lol no silly rabbit i meant \[7\div 282\] you can do \[.07\div 28.8\] but that is the same number doe

OpenStudy (rocketisawesome):

there

OpenStudy (rocketisawesome):

that is how you do it

OpenStudy (rocketisawesome):

Your welcome

OpenStudy (sasquat):

uh i dont think so cuz 2.83 is not ur base

OpenStudy (rocketisawesome):

hold on sorry i messed up the thing

OpenStudy (rocketisawesome):

i meant

OpenStudy (rocketisawesome):

2.82 not 2.83

OpenStudy (danieke):

the answer was 2.5

OpenStudy (rocketisawesome):

no the answer is 2.12%

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new~number - old~number}{old~number } \times 100\)

OpenStudy (rocketisawesome):

That's what bI did mathstudent that is what I showed them

OpenStudy (rocketisawesome):

I did that

OpenStudy (sasquat):

ok u know what i think misty is right, sorry danieke i messed up

OpenStudy (rocketisawesome):

People people. I am actually correct. Because of the fact that I actually did this problem and got it correct.

OpenStudy (mathstudent55):

\(percent~change = \dfrac{2.89 - 2.82}{2.82} \times 100 = 2.5\%\)

OpenStudy (rocketisawesome):

actually

OpenStudy (rocketisawesome):

you know what lets just give mathstudent the medal

OpenStudy (mathstudent55):

Since the percent change is positive, it is a percent increase.

OpenStudy (rocketisawesome):

the answer is 2.5 i did the math wrong i did 2.83

OpenStudy (rocketisawesome):

instead of 2.82

OpenStudy (sasquat):

and i totally did the math wrong

OpenStudy (mathstudent55):

You can memorize the formula for percent change, but it's important to understand it. The first step is to subtract the original price from the new price. That gives you the change. In this case, 2.89 - 2.82 = 0.07 The change in price was $0.07

OpenStudy (rocketisawesome):

5 people are viewing this

OpenStudy (rocketisawesome):

THE FINAL ANSWER IS.

OpenStudy (rocketisawesome):

THE PERCENT INCREASE IS 2.5% PERCENT

OpenStudy (sasquat):

actually 4 are viewing

OpenStudy (mathstudent55):

Now we want to know what percent of the original price is the change. 0.07 is what percent of 2.82? To find that, we divide the change by the ORIGINAL price, and multiply by 100. \(\dfrac{0.07}{2.82} \times 100 = 2.482...\%\) Since the problem asks to round to the nearest tenth of a percent, the answer is 2.5%

OpenStudy (danieke):

Thank you all but i have more questions

OpenStudy (rocketisawesome):

Ok

OpenStudy (rocketisawesome):

Sure what else?